OpenStudy (anonymous):

solve the equation by factoring x squared - 2x - 15 = 0 x squared - 11x + 30 = 0 x squared - x -12= 0 I need help on these three can anybody help with the steps ?

OpenStudy (tkhunny):

You need factors of the constant term from which you can build the middle term 15 = 3*5 -- See what you can get from that. 30 = 6*5 12 = 3*4 That's about all you need.